OpenStudy (anonymous):

N=-2x^2 +76x 430 How would I rewrite this equation with the right side completely factored

OpenStudy (ash2326):

N=-2x^2+76x+430 is this the equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you see the first question that i posted?

OpenStudy (anonymous):

ash2326 are you there?

OpenStudy (anonymous):

try quadratic formula maybe

OpenStudy (ash2326):

sorry, i was busy, wait I'm solving

OpenStudy (anonymous):

ok thank you!

OpenStudy (ash2326):

N=-2x^2+76x+430 ax^2+bx+c=0 x=(-b+root(b^2-4ac))/2a x=(-b-root(b^2-4ac))/2a a=-2 b=76 c=430 x=(-76+root(76^2+4*2*430)/-4 x=(-76+root(5776+3440)/-4 x=(-76+root(9216))/-4 x=(-76+96)/-4 x=(-76-96)/-4 so x=-5 and x=43 N= -2*(x+5)(x-43)

OpenStudy (anonymous):

x = 43 or x = -5

OpenStudy (anonymous):

using quadratic formula